There was a problem with Kohzu stage for 1064 crystals of Fe HRM. It does not move equidistantly. Problem seems to be solved by release of 2 (of 4) screws fixing motors on the stage.For details see figure in NRS setup and control

SmarAct rotation stage

Hendrik lindemann (4425) took one smaract rotation stage with sensormodule and cabel.

LEMO 50 Ohm

Today I took the only LEMO 50 Ohm resistor I could find.

Now both LEMO boxes (EH2 + EH3) have none left.

Is there another stock or do we need to order some?!?

Lemo cabel test sys

Mario Wendt (3115) has taken the Lemo cabel test sys. for this week.

The beamposition of the beam was in random intervals (order of 1s) jumping in y direction (different amount, order about 1mm at beginning of EH2).

We finally switched off the fuse of the dcm piezo and the problem was solved.

Undulator didn't move

When we moved Undulator A to its nominal minimum gap of 12.7mm, it stopped at 12.8mm and could not be moved anymore.

Neither init nor reset motor from the ATK panel worked. The machine could not move the undulator, either.

The shift service suggested to call Andreas Schoeps, who checked from home and found the problem:

  • The undulator had not reached its "normal" limit switch, but was stopped by the "emergency" limit switch, which prevents the undulator from touching the vacuum chamber.
  • Andreas set the minimum limit for Undulator A to 12.9mm, which is safe now.
  • He thinks that either the vacuum chamber and the undulator have moved relative to each other, or the mounting of the limit switch might have mechanical problems.

In winter shutdown this should be checked by undulator group.

Dear all,

I couldn't sit back and see the misery any more.....and ordered new torches!

Today I replaced the old  glooming dim-light torch-crab in the tool-trolleys by brandnew rechargeable LED torches. You can either charge them via USB or wall plug.

Enjoy them and....feel "enlightened" (smile)



Broken X 95 clamps

Dear All,

whenever you come across sluggish, stiff or broken x 95 clamps, please

DON'T just put them back into the carbinet and take a good one.

Please put them in the NEW BOX IN THE CABINET shown on the picture below.

I will repair them. Thanks !

Macro mv2max is corrected.

DESY standard library was changed ~ 2 weeks ago (function "HasyUtils.createScanInfo" was changed). As result, our P01 macro "mv2max" did not work anymore.

Now, macro "mv2max" is corrected and work again. If it does not work, please, reconfig sardana.

Dear all,

the K2Fe(CN)6 samples and a new 57Fe foil arrived. I put the Fe foil together with others Fe samples and the 119 Sn foil into a new box which replaces the small white box we had. The potasssium magnesium ferrocyanide came with its own box, see photo.

The boxes are stored in the dry cabinet in CH1.

Here are the Mössbauer data sheets of the Fe foil and the K2Fe(CN)6 samples. They can also be found on paper in a folder in CH1.

and put to the desk in EH1. Hans-Christian checked influence on measurements in EH3 - everything O.K.

Beamline Laptop

currently it is in the cabinet in Room L116 (office of Conrad, Simon, Frank-Uwe).

Please leave a message here, when you take it!

Frank-Uwe: 08.11.17 Beamline Laptop is in OH2 / 10.11.17 it is back to office

Frank-Uwe: 14.11.17 Laptop is in EH2

Frank-Uwe: 15.11.17 MaxV Card changed. No other changes done.

On Sunday evening we experienced a problem with moving motors of 2nd artificial channel cut  hrm1064.

  • Some of the motors did not move anymore, which happened during a long scan.
  • It concerned motors of hrm1064 on ZMX-crate 128 in EH1, slots 33 to 36.
  • We tried ZMX-card 40 in the same crate, motor hrm1064_x did not move.
  • The motor connected to ZMX-card 41 in the same crate worked.
  • We concluded that the first 8 ZMX-cards, which are controlled by the same OMSVME have the problem, means probably this VME-card is the reason.
  • We shifted the four hrm1064 motors by 8 slots:
    hrm1064_tilt  33 →  41
    hrm1064_th   34 →  42
    hrm1064_y    35 →  43
    hrm1064_x    36 →  44
  • Correspondingly the online.xml was modified

 After that operation the hrm1064 motors can be controlled again from spock.

There was no possibility to control Vacuum Interlock system due to the server ERROR.

One still can see the system via: